# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 4. Traditional Adversarial Attacks
We divide traditional adversarial attacks into two stages: the basic strategy (Stage 1) and attack enhancement (Stage 2). In the basic strategy stage, researchers explore and adapt common problem-solving approaches from other fields for use in adversarial attacks, thereby developing a foundational framework. Methods from this stage often serve as a basis for future approaches. In the attack enhancement stage, the design of attack methods typically follows specific motivations. For instance, these methods aim to generate AEs under constraints such as limited or no access to the victim model, or to improve the stealthiness, physical robustness, and generation speed of the AEs.
@@
4.1. Stage 1: The Basic Strategy
As shown in Fig. 3, attack methods at this stage can be categorized into five types. Single-step and iterative methods typically generate AEs by adding gradients that differ from the true prediction to the original sample. Optimization-based methods approach the generation of perturbations as an optimization problem, while search-based and generative-model-based methods utilize search algorithms or rely on generator to generate AEs.
@@ 4.1. Stage 1: The Basic Strategy
Optimization-based methods (Carlini and Wagner, 2017; Szegedy, 2013; Biggio et al., 2013) transform the box constraint (Carlini and Wagner, 2017) on perturbations into an optimization objective (such as the P-norm) and can use algorithms like Adam (Kingma, 2014) to generate adversarial perturbations. Similar to iterative methods, these approaches can create detailed perturbations to enhance stealthiness but at the cost of transferability (Dong et al., 2019; Lin et al., 2019) and generation speed (Poursaeed et al., 2018; Moosavi-Dezfooli et al., 2016; Li et al., 2020b; Xie et al., 2019b).
Search-based methods can be divided into two types: heuristic and custom search methods. Heuristic search methods (Nguyen et al., 2015; Sharif et al., 2016; Su et al., 2019; Tseng et al., 2024) generate AEs by relying solely on evaluative information like fitness (e.g., confidence scores), often obtained through queries to the victim model, making them inherently gray-box methods. Custom search methods assist attacks by identifying decision boundaries (Brendel et al., 2017) and vulnerable pixel locations (Papernot et al., 2016b; Narodytska and Kasiviswanathan, 2016). These methods may modify only a few pixels (Papernot et al., 2016b; Su et al., 2019; Narodytska and Kasiviswanathan, 2016) or regions (Sharif et al., 2016), providing a degree of stealth. However, high query (Su et al., 2019; Narodytska and Kasiviswanathan, 2016; Tseng et al., 2024; Brendel et al., 2017) or computation (Papernot et al., 2016b) counts limit their practical application.
Generating AEs using generative models (Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018; Zhou et al., 2023b, a; Chen et al., 2023a; Guo et al., 2024) has two key pros: 1) rapid generation speed; and 2) high sample naturalness. In common-used generators, autoencoders (Baluja and Fischer, 2017; Poursaeed et al., 2018) and GANs (Xiao et al., 2018; Zhou et al., 2023b, a) can produce perturbations in a single forward process, significantly enhancing generation speed. Although diffusion models require iterative denoising, they also maintain relatively fast speeds (Chen et al., 2023a; Guo et al., 2024). This method often avoids using box constraints to limit perturbation size and instead aims to create perceptually invisible AEs (Unrestricted AEs, UAEs (Chen et al., 2023a)), redefining the concept of stealthiness for AEs from a fresh perspective.
4.2. Stage 2: Attack Enhancement
As illustrated in Fig. 3, attacks at this stage can be categorized into three groups based on the attacker’s level of knowledge. Additionally, methods from Stage 1 may reappear here as seminal works viewed from a different perspective.
@@ 4.2.1. White-box
In the white-box scenario, generating AEs serves primarily two purposes: 1) evaluating model robustness and 2) enhancing specific attributes of samples, such as physical robustness, generation speed, and stealthiness. As noted by Carlini and Wagner (Carlini and Wagner, 2017), only sufficiently strong AEs can accurately measure the true lower bound of model behavior, representing the upper bound of robustness under attack. Consequently, white-box attack methods aimed at robustness evaluation focus on enhancing adversariality. For instance, FAB (Croce and Hein, 2020a) generates attack samples close to the decision boundary by iteratively linearizing the classifier and projecting. CW (Carlini and Wagner, 2017) and PGD (Madry, 2017) implement effective attacks using optimization and iterative methods, respectively. Building on PGD, APGD (Croce and Hein, 2020b) improves the iterative process by dynamically adjusting the step size, while MT (Gowal et al., 2019) enhances diversity in starting points by maximizing changes in the output domain. Additionally, AA (Croce and Hein, 2020b) and CAA (Mao et al., 2021) seek to bolster evaluation capabilities by aggregating multiple attacks, whereas A3 (Liu et al., 2022) optimizes the attack process dynamically through adaptive adjustments of starting points and automatic selection of attack images.
Beyond adversariality, enhancing the performance of AEs in physical robustness, stealthiness, and generation speed is crucial. This topic will be addressed in section 5. While improvements in these capabilities are relevant for black-box and gray-box scenarios as well, the focus here is on white-box research due to its extensive exploration.
@@ 4.2.2. Gray-box
In the gray-box scenario, the attacker has access to limited information through queries, such as predicted labels (Narodytska and Kasiviswanathan, 2016; Brendel et al., 2017; Papernot et al., 2017, 2016a; Ilyas et al., 2018), label rankings (Ilyas et al., 2018), and confidence scores (Narodytska and Kasiviswanathan, 2016; Ilyas et al., 2018; Chen et al., 2017; Su et al., 2019; Tseng et al., 2024; Sharif et al., 2016; Nguyen et al., 2015). This limitation naturally gives rise to two attack strategies: 1) directly generating attack samples using the available information, and 2) estimating additional information, such as gradients (Chen et al., 2017; Ilyas et al., 2018; Narodytska and Kasiviswanathan, 2016), decision boundaries (Brendel et al., 2017), and model behavior (Papernot et al., 2017, 2016a), based on the limited data provided.
Direct generation methods often utilize heuristic search algorithms, as they require only fitness-related information from the victim model to evaluate the quality of individuals. Sample updates are guided by specific strategies that use fitness metrics to retain the best individuals, iterating ultimately to produce effective attack samples. EA-CPPN (Nguyen et al., 2015) and OPA (Su et al., 2019) generate AEs using genetic algorithms (GAs), a type of evolutionary algorithm, while RSA-FR (Sharif et al., 2016) and AE-QTS (Tseng et al., 2024) employ particle swarm and quantum-inspired algorithms, respectively, to implement their attacks.
In estimation-based methods, the estimated targets may include gradients, decision boundaries, and model behavior. ZOO (Chen et al., 2017) employs symmetric difference quotients to estimate gradients on a pixel-wise basis, followed by CW attacks. NES (Ilyas et al., 2018) estimates the gradient at the current iteration by using finite differences over Gaussian bases to update perturbations with PGD. In contrast to ZOO and NES, LSA (Narodytska and Kasiviswanathan, 2016) implicitly estimates the gradient saliency map (Simonyan, 2013) for pixel positions relative to true labels through a local greedy search, effectively reducing perturbation size by targeting the most sensitive areas for label prediction. BA (Brendel et al., 2017) uses rejection sampling to allow perturbed samples to randomly walk toward the decision boundary, indirectly estimating the victim model’s decision boundary. JDA (Papernot et al., 2017) and JDA+ (Papernot et al., 2016a) label a surrogate dataset by querying the victim model, train a surrogate model to mimic the victim’s behavior, and ultimately generate AEs using white-box methods on the surrogate model.
@@ 5. Motivations for Improving Adversarial Attacks
[(Dong et al., 2018; Lin et al., 2019; Wang and He, 2021; Gao et al., 2020b; Huang et al., 2019; Gao et al., 2020a; Li et al., 2020c, b; Qin et al., 2022; Chen et al., 2023c) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Qin et al., 2022; Chen et al., 2023c) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Zhou et al., 2018; Sharma et al., 2018; Dong et al., 2019; Li et al., 2020a; Gao et al., 2020b, a; Sharma et al., 2019) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Guo et al., 2024; Poursaeed et al., 2018; Zhou et al., 2023b, a; Moosavi-Dezfooli et al., 2017; Liu et al., 2016; Xiong et al., 2022; Huang et al., 2023a; Xie et al., 2019b; Dong et al., 2019; Lin et al., 2019; Xie et al., 2017b; Yan et al., 2022; Long et al., 2022; Wang et al., 2021) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Papernot et al., 2016b; Su et al., 2019; Narodytska and Kasiviswanathan, 2016) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Chen et al., 2023a; Guo et al., 2024; Liu et al., 2023c; Laidlaw et al., 2020; Zhao et al., 2020; Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
Speed (section 5.4), color=majorblue, fill=majorblue!15, text width=2.5cm, text=black
@@
[(Zhang et al., 2020; Pintor et al., 2021; Croce and Hein, 2020a; Wu et al., 2020) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018; Zhou et al., 2023b, a; Chen et al., 2023a; Guo et al., 2024) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
Fig. 4 depicts the motivations for enhancing the capabilities of AEs, which were omitted in section 4.2. These include improving transferability (section 5.1), physical robustness (section 5.2), stealthiness (section 5.3), as well as boosting generation speed (section 5.4).
@@ 5.1.1. Iteration Optimizing
MI-FGSM (Dong et al., 2018), NI-FGSM (Lin et al., 2019), and VI-FGSM (Wang and He, 2021) enhance transferability based on BIM (Kurakin et al., 2018) by incorporating momentum, applying Nesterov gradient descent for better step prediction, and tuning gradient variance between perturbed points and their neighbors, respectively. Both PI-FGSM (Gao et al., 2020b) and ILA (Huang et al., 2019) follow the principle that greater perturbation norm leads to greater transferability. PI-FGSM diffuses excessive gradients through smoothing convolution, while ILA increases the perturbation norm on features. PI-FGSM++ (Gao et al., 2020a) and ILA++ (Li et al., 2020c) further improve transferability by incorporating a temperature term during model aggregation and utilizing intermediate results from different time steps. POM (Li et al., 2020b) introduces the Poincaré distance as a similarity metric, addressing noise decay in targeted transfer attacks, where shrinking gradients over iterations reduce perturbation diversity and adaptability. To mitigate overfitting to surrogate models, RAP (Qin et al., 2022) and CWA (Chen et al., 2023c) aim to make the perturbed sample converge in flatter regions of the loss landscape, where the transferability is shown to be higher (Wu et al., 2017; Li et al., 2018; Chen et al., 2022).
@@ 5.1.2. Flat Area Encouraging
Previous studies (Wu et al., 2017; Li et al., 2018; Chen et al., 2022) have shown that a smaller Hessian matrix norm indicates flatter regions in the objective function, which correlates with better generalization. However, the high computational cost limits the use of the Hessian norm. To address this, both CWA (Chen et al., 2023c) and RAP (Qin et al., 2022) adopt a Min-MAX bi-level optimization approach to find flatter regions. CWA employs Sharpness-Aware Minimization (Foret et al., 2020), alternating between gradient ascent and descent steps to promote convergence to flat regions. RAP, in contrast, searches for local points with higher loss in the inner loop and minimizes the loss at these points in the outer loop, ensuring that the neighboring areas around the perturbed sample also have lower loss values, thus achieving convergence in flatter regions.
@@ 5.1.3. Smooth Encouraging
Another concept related to smoothing is aggregation. Aggregation on pixels refers to making perturbations exhibit regional homogeneity (RH). Li et al. (Li et al., 2020a) found that perturbations generated using adversarial or defensive models as surrogates tend to have coarser granularity and exhibit certain structural patterns. They termed this feature RH and argued that it helps improve transferability in black-box attacks. RHP (Li et al., 2020a) trains a transformation module on AEs to convert normal perturbations into ones with RH. During training, the transformation module is encouraged to underfit, thereby implicitly generating universal perturbations. Inspired by RHP (Li et al., 2020a) and Rosen’s gradient projection (Rosen, 1960), PI-FGSM (Gao et al., 2020b) disperses exceeding perturbations to surrounding areas using a uniform kernel, updating perturbations for a local patch to enhance aggregation. PI-FGSM++ (Gao et al., 2020a) further enhances targeted transferability by adding a temperature term when aggregating models’ logits, which requiring higher confidence for misclassification.
Both smoothing and aggregation are ways to achieve low-frequency perturbations (Sharma et al., 2019). (Wang et al., 2020a, b; Yin et al., 2019) explained that the adversarial nature of a sample may come from high-frequency signal interference, while low-frequency perturbations have been found to improve the transferability (Sharma et al., 2019; Zhou et al., 2018; Sharma et al., 2018; Dong et al., 2019; Li et al., 2020a; Gao et al., 2020b, a) and physical robustness (Thys et al., 2019; Komkov and Petiushko, 2021; Sharif et al., 2016). For physical robustness, these methods introduce TV loss (Rudin et al., 1992) to smooth the perturbations, making them more stable under image interpolation operations across different devices and less visible to the human eye (Komkov and Petiushko, 2021).

### v2

@@ 4. Traditional Adversarial Attacks
Researchers often begin by applying established methods to new problems before developing enhancements based on specific motivations. Inspired by this, we identify two key attributes of adversarial attacks: Basic Generation Methods (classical AE generation strategies, Attribute 1) and Enhancement Motivations (motivational drivers under different knowledge constraints, Attribute 2). These dimensions represent distinct yet interconnected facets of attack design.
4.1. Attribute 1: Basic Generation Methods
As shown in Fig. 4, attacks under this attribute can be categorized into five types. Single-step and iterative methods typically generate AEs by adding gradients from varying malicious purposes (e.g., optimizing a spiteful loss). Optimization-based methods approach the generation of perturbations as an optimization problem, while search-based and generative-model-based methods utilize search algorithms or rely on generator to generate AEs.
@@ 4.1. Attribute 1: Basic Generation Methods
Optimization-based methods (Carlini and Wagner, 2017; Szegedy et al., 2014; Biggio et al., 2013) convert box constraints on perturbations into optimization objectives, using algorithms (e.g., Adam (Kingma, 2014)) to generate AEs. Similar to iterative methods, they produce precise perturbations for enhanced stealthiness but compromise transferability (Dong et al., 2019; Lin et al., 2020) and efficiency (Poursaeed et al., 2018; Moosavi-Dezfooli et al., 2016; Li et al., 2020a; Xie et al., 2019b).
Search-based methods can be divided into two types: heuristic and custom search methods. Heuristic search methods (Nguyen et al., 2015; Sharif et al., 2016; Su et al., 2019; Tseng et al., 2024) generate AEs by relying solely on evaluative information like fitness (e.g., confidence scores), often obtained through queries to the victim model, making them inherently gray-box methods. Custom search methods assist attacks by identifying decision boundaries (Brendel et al., 2017) and vulnerable pixel locations (Papernot et al., 2016c; Narodytska and Kasiviswanathan, 2016). These methods may modify only a few pixels (Papernot et al., 2016c; Su et al., 2019; Narodytska and Kasiviswanathan, 2016) or regions (Sharif et al., 2016), providing a degree of stealth. However, high query (Su et al., 2019; Narodytska and Kasiviswanathan, 2016; Tseng et al., 2024; Brendel et al., 2017) or computation (Papernot et al., 2016c) counts limit their practical application.
Generating AEs using generative models (Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018; Zhou et al., 2023b, a; Chen et al., 2023b; Guo et al., 2024) has two key pros: 1) rapid generation speed; and 2) high sample naturalness. In commonly-used generators, autoencoders (Baluja and Fischer, 2017; Poursaeed et al., 2018) and GANs (Xiao et al., 2018; Zhou et al., 2023b, a) can produce perturbations in a single forward process, significantly enhancing generation speed. Although diffusion models require iterative denoising, they also maintain relatively fast speeds (Chen et al., 2023b; Guo et al., 2024). This method often avoids using box constraints to limit perturbation and instead aims to create perceptually invisible AEs (Unrestricted AEs, UAEs (Chen et al., 2023b)), redefining the concept of stealthiness for AEs from a fresh perspective.
4.2. Attribute 2: Enhancement Motivations
As illustrated in Fig. 4, attacks under this attribute can be categorized into three groups based on the different knowledge level. Attribute taxonomy classifies attacks from different perspectives, allowing a single attack to fall under both techniques (Attribute 1) and motivations (Attribute 2).
@@ 4.2.1. White-box
In white-box scenarios, AEs serve primarily two purposes: 1) evaluating model robustness and 2) enhancing attributes, like physical robustness, generation efficiency, stealthiness, and interpretability. As noted by Carlini and Wagner (Carlini and Wagner, 2017), only highly potent AEs can accurately assess the true lower bound of model behavior, representing the upper robustness limit under attack. Consequently, white-box attacks for robustness evaluation focus on maximizing adversariality. Specifically, FAB (Croce and Hein, 2020) generates boundary-proximate AEs via iterative classifier linearization and projection, while CW (Carlini and Wagner, 2017) and PGD (Madry, 2017) employ optimization and iterative techniques, respectively. APGD (Croce et al., 2020) refines PGD by dynamically adjusting step size, and MT (Gowal et al., 2019) by enhancing the diversity of starting points. On attack automation, AA (Croce et al., 2020) and CAA (Mao et al., 2021) aggregate multiple attacks, whereas A3 (Liu et al., 2022) try to adaptively adjust starting points and automatically select attack images. (Agnihotri et al., 2024; Gu et al., 2022) adapts to dense prediction tasks (e.g., segmentation) by dynamically adjusting perturbation weights of pixels. Beyond adversariality, improving AEs in physical robustness, stealthiness, generation efficiency, and interpretability remains critical (see section 5). While relevant to black-box and gray-box settings, this discussion focuses on white-box research due to its extensive study.
@@ 4.2.2. Gray-box
In gray-box scenarios, attackers access limited information via queries, including predicted labels (Narodytska and Kasiviswanathan, 2016; Brendel et al., 2017; Papernot et al., 2017, 2016b; Ilyas et al., 2018), label rankings (Ilyas et al., 2018), and confidence scores (Narodytska and Kasiviswanathan, 2016; Ilyas et al., 2018; Chen et al., 2017; Su et al., 2019; Tseng et al., 2024; Sharif et al., 2016; Nguyen et al., 2015; Li et al., 2024d). This limitation prompts two strategies: 1) crafting AEs directly from available data, or 2) inferring additional information like gradients (Chen et al., 2017; Ilyas et al., 2018; Narodytska and Kasiviswanathan, 2016), decision boundaries (Brendel et al., 2017), or model behavior (Papernot et al., 2017, 2016b) from the limited data provided.
Direct generation methods often utilize heuristic search algorithms, relying solely on fitness-related information from victim models to evaluate individuals. These approaches iteratively refine solutions by retaining high-fitness individuals, ultimately producing effective AEs. EA-CPPN (Nguyen et al., 2015) and OPA (Su et al., 2019) use genetic algorithms (GAs), while RSA-FR (Sharif et al., 2016) and AE-QTS (Tseng et al., 2024) employ PSO and quantum-inspired algorithms to generate AEs. Optimized by differential evolution, (Li et al., 2023c, 2022b) construct searchable weights based on image differences to reduce the dimensionality of the search space, enhancing search efficiency. Recently, FTQ-PSO (Li et al., 2024d) integrated transfer attacks with PSO. It first leveraging a GAN to generate an initial perturbation, then refining it by adjusting the low-dimensional input of the GAN with PSO.
In estimation-based methods, the estimated targets may include gradients, decision boundaries, and model behavior. ZOO (Chen et al., 2017) employs symmetric difference quotients to estimate gradients on a pixel-wise basis, followed by CW attacks. NES (Ilyas et al., 2018) estimates the gradient at the current iteration by using finite differences over Gaussian bases to update perturbations with PGD. In contrast to ZOO and NES, LSA (Narodytska and Kasiviswanathan, 2016) implicitly estimates the gradient saliency map (Simonyan, 2013) for pixel positions relative to true labels through a local greedy search, effectively reducing perturbation size by targeting the most sensitive areas for label prediction. BA (Brendel et al., 2017) uses rejection sampling to allow perturbed samples to randomly walk toward the decision boundary, indirectly estimating the victim model’s decision boundary. JDA (Papernot et al., 2017) and JDA+ (Papernot et al., 2016b) label a surrogate dataset by querying the victim model, train a surrogate model to mimic the victim’s behavior, and ultimately generate AEs using white-box methods on the surrogate model.
@@ 5. Motivations for Improving Adversarial Attacks
[(Dong et al., 2018; Lin et al., 2020; Wang and He, 2021; Gao et al., 2020b; Huang et al., 2019; Gao et al., 2020a; Li et al., 2020c, a; Qin et al., 2022; Chen et al., 2023d) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Qin et al., 2022; Chen et al., 2023d) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Zhou et al., 2018; Sharma et al., 2018; Dong et al., 2019; Li et al., 2020b; Gao et al., 2020b, a; Sharma et al., 2019) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Guo et al., 2024; Poursaeed et al., 2018; Zhou et al., 2023b, a; Moosavi-Dezfooli et al., 2017; Liu et al., 2016; Xiong et al., 2022; Huang et al., 2023a; Xie et al., 2019b; Dong et al., 2019; Lin et al., 2020; Xie et al., 2017b; Yan et al., 2022; Long et al., 2022; Wang et al., 2021c) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Papernot et al., 2016c; Su et al., 2019; Narodytska and Kasiviswanathan, 2016) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Chen et al., 2023b; Guo et al., 2024; Liu et al., 2023c; Laidlaw et al., 2020; Zhao et al., 2020; Laidlaw and Feizi, 2019; Shamsabadi et al., 2021, 2020) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
Efficiency (section 5.4), color=majorblue, fill=majorblue!15, text width=2.5cm, text=black
@@
[(Zhang et al., 2020; Pintor et al., 2021; Croce and Hein, 2020; Wu et al., 2020) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
[(Baluja and Fischer, 2017; Poursaeed et al., 2018; Xiao et al., 2018; Zhou et al., 2023b, a; Chen et al., 2023b; Guo et al., 2024) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
Interpretability (section 5.5), color=majorblue, fill=majorblue!15, text width=2.5cm, text=black
Output-Interpretation Mismatch, color=majorblue, fill=majorblue!15, text width=4.5cm, text=black
[(Vadillo et al., 2025) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
Output-Interpretation Match, color=majorblue, fill=majorblue!15, text width=4.5cm, text=black
[(Vadillo et al., 2025) , color=majorblue, fill=majorblue!15, text width=3.6cm, text=black]
@@
Fig. 5 depicts the motivations to enhance AEs, which were omitted in section 4.2. These include improving transferability (section 5.1), physical robustness (section 5.2), stealthiness (section 5.3), generation efficiency (section 5.4), and interpretability of results (section 5.5).
@@ 5.1.1. Iteration Optimizing
MI-FGSM (Dong et al., 2018), NI-FGSM (Lin et al., 2020), and VI-FGSM (Wang and He, 2021) enhance transferability based on BIM (Kurakin et al., 2018) by incorporating momentum, applying Nesterov gradient descent for better step prediction, and tuning gradient variance between perturbed points and their neighbors, respectively. Both PI-FGSM (Gao et al., 2020b) and ILA (Huang et al., 2019) follow the principle that greater perturbation norm leads to greater transferability. PI-FGSM diffuses excessive gradients through smoothing convolution, while ILA increases the perturbation norm on features. PI-FGSM++ (Gao et al., 2020a) and ILA++ (Li et al., 2020c) further improve transferability by absorbing a temperature term during aggregation and utilizing intermediate results from multiple time steps. On feature domains, FIA (Wang et al., 2021b) and NAA (Zhang et al., 2022b) enhance transferability by suppressing positive and amplifying negative features. POM (Li et al., 2020a) introduces the Poincaré distance as a similarity metric, addressing noise decay in targeted transfer attacks, where shrinking gradients over iterations reduce perturbation diversity and adaptability.
@@ 5.1.2. Flat Area Encouraging
Previous studies (Wu et al., 2017; Li et al., 2018; Chen et al., 2022) have shown that a smaller Hessian matrix norm indicates flatter regions in the objective function, which correlates with better generalization. However, the high computational cost limits the use of the Hessian norm. To address this, both CWA (Chen et al., 2023d) and RAP (Qin et al., 2022) adopt a Min-MAX bi-level optimization approach to find flatter regions. CWA employs Sharpness-Aware Minimization (Foret et al., 2020), alternating between gradient ascent and descent steps to promote convergence to flat regions. RAP, in contrast, searches for local points with higher loss in the inner loop and minimizes the loss at these points in the outer loop, ensuring that the neighboring areas around the perturbed sample also have lower loss values, thus achieving convergence in flatter regions.
@@ 5.1.3. Smooth Encouraging
Another concept related to smoothing is aggregation. Aggregation on pixels refers to making perturbations exhibit regional homogeneity (RH). Li et al. (Li et al., 2020b) found that perturbations generated using adversarial or defensive models as surrogates tend to have coarser granularity and exhibit certain structural patterns. They termed this feature RH and argued that it helps improve transferability in black-box attacks. RHP (Li et al., 2020b) trains a transformation module on AEs to convert normal perturbations into ones with RH. During training, the transformation module is encouraged to underfit, thereby implicitly generating universal perturbations. Inspired by RHP (Li et al., 2020b) and Rosen’s gradient projection (Rosen, 1960), PI-FGSM (Gao et al., 2020b) disperses exceeding perturbations to surrounding areas using a uniform kernel, updating perturbations for a local patch to enhance aggregation. PI-FGSM++ (Gao et al., 2020a) further enhances targeted transferability by adding a temperature term when aggregating models’ logits, which requiring higher confidence for misclassification.
Both smoothing and aggregation are ways to achieve low-frequency perturbations (Sharma et al., 2019). (Wang et al., 2020a, b; Yin et al., 2019) explained that the adversarial nature of a sample may come from high-frequency signal interference, while low-frequency perturbations have been found to improve the transferability (Sharma et al., 2019; Zhou et al., 2018; Sharma et al., 2018; Dong et al., 2019; Li et al., 2020b; Gao et al., 2020b, a) and physical robustness (Thys et al., 2019; Komkov and Petiushko, 2021; Sharif et al., 2016). For physical robustness, these methods introduce TV loss (Rudin et al., 1992) to smooth the perturbations, making them more stable under image interpolation operations across different devices and less visible to the human eye (Komkov and Petiushko, 2021).